OpenStudy (anonymous):

Find the equation for the line through (2.6) and perpendicular to y=-5/4x+1 .

OpenStudy (jackellyn):

The slope of the other line that is perpendicular is the negated reciprocal so -5/4=4/5 for new line using the formula y-y1=m(x-x1) plug in values y-6=4/5(x-2) y-6=4/5x + (4/5)(-2) y= 4/5x -(4/5*2)+6 y= 4/5x + 22/5

OpenStudy (anonymous):

Perpendicular, so the slope is the opposite inverse of -5/4 So slope=4/5 Now we use this: y-y1=m*(x-x1) and plug x1=2, y1=6 So then: y-6=4/5*(x-2) y=4/5x-8/5+6 y=4/5x+22/6

OpenStudy (anonymous):

Sorry, y=4/5x+22/5